40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1478     Entry time: Mon Apr 13 17:55:37 2009
Author: Jenne 
Type: Update 
Category: PSL 
Subject: PMC LO Mon Calibration 

I have calibrated the PMC LO Mon (C1:PSL-PMC_LODET) on the PMC's EPICS screen, by inputting different RF LO levels into the LO input of the PMC servo board.

 

Since the RF output adjust slider on the PMC's Phase Shifter screen doesn't do a whole lot (see elog 1471), I used a combination of attenuators and the slider to achieve different LO levels. I measured the level of the attenuated RF out of the LO board using the 4395A in spectrum analyzer mode, with the units in dBm, with 50dB attenuation to make it stop complaining about being overloaded.  For each row in the table I measured the RF level using the 4395, then plugged the cable back into the PMC servo board to get the EPICS screen's reading.

The last 2 columns of the table below are the 'settings' I used to get the given RF LO level. 

RF LO Input to PMC Servo Board [dBm] LO Mon on EPICS Screen [no units] RF Output Adjust Slider [V] Attenuators used [dB]
16.004 +- 0.008 0.1200 +- 0.0003 0 0
15.001 +- 0.004 0.0708 +- 0.0008 0 1
14.079 +- 0.008 0.0318 +- 0.0001 8 1
13.002 +- 0.006 0.0126 +- 0.0004 0 3
11.992 +- 0.010 0.0024 +- 0.0008 0 4
10.994 +- 0.010 -0.0024 +- 0.0003 0 4+1=5
9.993 +- 0.008 -0.0047 +- 0.0007 0 3+3=6

 

When the new mixers that Steve ordered come in (tomorrow hopefully), I'll put in a Level 13 mixer in place of the current Level 23 mixer that we have.  Also, Rana suggested increasing the gain on the op-amp which is read out as the LO Mon so that 13dBm looks like 1V.  To do this, it looks like I'll need to increase the gain by ~80.  

Attachment 1: LOmonCalibration.png  8 kB  | Hide | Hide all
LOmonCalibration.png
ELOG V3.1.3-